A steam engine develops 600 kW at 180 r.p.m. The coefficient of fluctuation of energy is

found to be 0.5 and the fluctuation of speed is required to be kept within ±1% of the mean

speed. The maximum fluctuation of the energy during operation is given by:


1
Expert's answer
2021-03-09T04:54:46-0500

P = 600kW

"\omega" = 180 r.p.m = "\dfrac{180×2π}{60}" = 18.85 rad/s


fluctuation of speed = 1% × 18.85 = 0.1885


maximum flucctuation of energy = 600/0.1885 = 3181kJ